OpenStudy (anonymous):

2x2+6ax+a=0 When is the integer value A sum of the squares of the roots of the equation is 38

OpenStudy (anonymous):

some can decide this stand my best friend for eternity?

OpenStudy (goformit100):

Sure one Guy can Help You, He is @Dharmaputra_993 .

OpenStudy (anonymous):

\[2x ^{2}+6ax+a =0\] When is the integer value A sum of the squares of the roots of the equation is 38? I didn't understand the question, sorry.

OpenStudy (anonymous):

you just must find value A ? you have (x2+x2)=38

OpenStudy (anonymous):

Dharmapatura So ?

OpenStudy (anonymous):

Okay got it, so \[x ^{2}+x ^{2}=38\] \[2x ^{2}=38\] \[x ^{2}=38/2=19\] \[x=\sqrt{19}\] If that's what you wanted.

OpenStudy (anonymous):

not it such P2-2q viet theoremy you must use

OpenStudy (anonymous):

Can you take a screenshot of the question and put it like an image here please?

OpenStudy (anonymous):

ax2 +px+c=0 x1+x2=-p x1+x2=-c